Grade 11Mechanics

2kg block start from rest on the positive x axis 3m from origin and has acceleration given by a=(4m/s^2)i+ (3m/s^2)j.at the end of 2s its angular momentum about origin is

Rishav raj
7 Years agoGrade 11

1 Answer

Piyush Kumar Maurya
7 Years ago
Initially displacement vector Ro= 3i cap 
Velocity, V = a.t = 4t icap + 3t jcap
Final Displacement,Rf= ∫ V.dt = ∫(4t icap +3t jcap) dt
 
= 2t2 icap + 3t2/2 jcap
Displacement , vector R= Rf + Ro 
                                         = (3 + 2ticap +  3t2/2 jcap)
At t= 2s 
Displacement,vector R = 11 icap + 6 jcap
Velocity, V = 8 icap + 6 jcap
 
Vector L = vector R x vector P
               = m (vector R x vector V)
    = 2 [( 11 icap + 6 jcap) x ( 8 icap + 6 jcap)]
    = 2( 66 kcap - 48 kcap)
    = 2( 18 kcap)
Vector L   = 36 kcap